Between Ithaca and Geneva, NY at the old Seneca Army Depot is a herd of white tailed deer (Odocoileus virginianus) which are normally brown.  The base is closed, but offers tours to see the deer herds. There are 800 deer in the population, 200 of which are white (yy).

Typically expression of this recessive allele makes the individuals more susceptible to predation, but the old army base is fenced off, providing protection from most predators. These brown-eyed deer are not albinos, but rather leucistic (lacking pigmentation in their hair only).  Management of the Seneca white deer on the former depot increased the proportion of deer exhibiting the trait

a.)What is the frequency of the recessive (white) allele in the population?

b.  What is the predicted frequency of carriers of the leucistic trait in the population?  
Answer #1

Answer:

a).

The frequency of white phenotype = yy = 200/800 = 0.25

The frequency of white allele = y = SQRT of 0.25 = 0.5

b).

The frequency of dominant allele = Y = 1-0.5 = 0.5 (as p+q=1)

The predicted frequency of carriers of the leucistic trait in the population = 2Yy = 2*0.5*0.5 = 0.5 or 50%